In a small country, the demand curve is given as: Q=100-5P, supply curve: Q=3P-12, and the world price is $10.

What is the social surplus under free trade?

If the government impose a $2/unit tariff on the good, what is the deadweight loss?

Show the change in equilibrium and deadweight loss on a graph.
Answer #1

equation d = 100-5P Demand When pro d = 100 when 0=o = 20 By plotting this we get the domestic demand curre d = 31-12 Supplydemand curre. ES - 10.5 (20-10) (50) C.S = $250 & Consumes susplus under Producer surplus is the triangle a and area above thhovernment impose a f 2 unit Now, the domestie &1lo+2) = $12. consumers paid For P - $12 Od = 100-5(12) for p = $12 Os = 3(12Government Revenue - (12-10)/40-24) $327 Deadweight loss = 10.5) (12-16)/2448) +0.5)(12-1050-4) = 6+ lo = 316 20 Domesti DOL